# LiV2O4: Hund-Assisted Orbital-Selective Mottness

### v1

@@ Abstract
We show that the remarkably small Fermi-liquid coherence scale and large effective mass observed in LiV2O4 are due to the proximity of a Hund-assisted orbital-selective Mott state. Our work is based on an ab initio dynamical mean-field approach, combining several quantum impurity solvers to capture the physics from high to very low temperature. We find that the Hund coupling plays a crucial role in rearranging the orbital populations and in generating the heavy mass and low coherence scale. The latter is found to be approximately 1–2 Kelvin, even though the most correlated orbital is found to be significantly doped ($\sim 10\%$) away from half-filling. A flat quasiparticle band appears near the Fermi level as a result of the strong electronic correlations. Finally, we discuss our results in comparison to experiments.
@@
Figure 1: (a) DFT band structure of LVO for the four $a_{1g}$- and eight $e_{g}^{\pi}$-like bands around the Fermi level. The eigenvalues of the MLWFs perfectly match the DFT bands [6]. Color: orbital character. Inset: Brillouin zone with high-symmetry points. (b) DFT DOS and projections onto the MLWFs.
@@
We show that LVO is close to a Hund-assisted orbital-selective Mott state [19, 20, 21, 14, 5]. The inter-orbital Hund coupling $J$ (and therefore the $e_{g}^{\pi}$ bands) plays a central role. First, it induces a redistribution of the ($a_{1g}$, $e_{g}^{\pi}$) orbital populations per vanadium due to interaction from $(0.4,2\times 0.55)$ in DFT to $(0.9,2\times 0.3)$ in DMFT. Second, it is responsible for the heavy mass and low $T_{\mathrm{FL}}\,$ via the spin-blocking mechanism characteristic of Hund systems in which electrons can only hop between atomic configurations with maximum spin [22, 23, 24, 25]. LVO is shown to be close to an orbital-selective Mott state in which the $a_{1g}$ orbital is prevented to fully localize at $T=0$ due to inter-orbital hopping [26]. Remarkably, we obtain a low $T_{\mathrm{FL}}\,$ already at $\sim\!10\%$ doping of the $a_{1g}$ band, in contrast to the $\sim\!2\%$ doping of Ref. [12].
@@
Figure 2: (a) $\mathrm{Im}\Sigma(i\omega_{n})$ for the $a_{1g}$ orbital (circles) and $e_{g}^{\pi}$ orbitals (crosses) at various $T$ from DFT+DMFT using QMC, TN, and NRG. Inset: zoom at low frequencies. (b) Scattering rate and (c) spectral function at the Fermi level as a function of $T$. Results at $i0^{+}$ in (b)–(c) for QMC and TN were obtained by fitting a fourth-order polynomial to the lowest eight Matsubara points. Insets in (b)–(c) magnify the FL behavior.
@@
Crossover to a heavy FL—The emergence of the low-$T$ heavy FL is illustrated in Fig. 2(a). It shows the imaginary part of the self-energy for each orbital as a function of Matsubara frequency for several $T$. The $a_{1g}$ self-energy has a strong $T$ dependence at low frequency, with a sudden decrease below the onset temperature $T_{\mathrm{onset}}\,$$\sim 12$ K. This is a hallmark of strong correlations and corresponds to a sharp crossover from a high-$T$ incoherent regime with a large scattering rate to a low-$T$ coherent regime. This is further supported by Fig. 2(b), which shows the extrapolated zero-frequency value of the $a_{1g}$ scattering rate $-\mathrm{Im}\Sigma_{a_{1g}}(i0^{+})$ as a function of $T$. The local spectral function (DOS) at the Fermi level $A_{a_{1g}}(0)$ correspondingly increases, as shown in Fig. 2(c). By contrast, the $e_{g}^{\pi}$ self-energy is smaller and depends more weakly on temperature and frequency.
The FL fully develops only below the FL coherence temperature $T_{\mathrm{FL}}\,$$\sim$ 1–2 K, an order of magnitude smaller than $T_{\mathrm{onset}}\,$. The FL regime is characterized by $-\mathrm{Im}\Sigma\!\sim\!\omega^{2}\!+\!\pi^{2}T^{2}$ (see End Matter for $T\!=\!0$ and Fig. 2(b) for $\omega\!=\!0$) for the retarded self-energy and $\chi^{\prime\prime}\!\sim\!\omega$ for the imaginary part of the retarded local spin susceptibility (see End Matter). The quasiparticle weights (from NRG at $T\!=\!0$) are $Z_{a_{1g}}\!\approx\!0.003$ and $Z_{e_{g}^{\pi}}\!\approx\!0.03$.
@@
To analyze the possible proximity of a Mott-insulating state, we perform a numerical experiment in which we add a crystal field to the Wannier Hamiltonian, while keeping the total $d$-electron count equal to 1.5 by adjusting the chemical potential. Figure 4(d) shows $T_{\mathrm{FL}}\,$determined by NRG as a function of the $a_{1g}$ orbital occupancy (lower horizontal scale) or the applied crystal field (upper scale). We see that $T_{\mathrm{FL}}\,$is driven to remarkably low values as half-filling of the $a_{1g}$ orbital is approached, with values as low as $10^{-6}$ K at $1\%$ doping. This can be interpreted as the system reaching an orbital-selective Mott (OSM) phase [19, 20, 21, 14, 5] in which the $e_{g}^{\pi}$ electrons remain itinerant while the $a_{1g}$ ones basically localize, full localization being prevented at $T\!=\!0$ by inter-orbital hopping [26]. The relevance of the OSM phenomenon and of Hund physics as a possible cause for $d$-electron HF behavior was recently emphasized by Crispino et al. [5] and also previously mentioned qualitatively for LVO in the NMR study of Shimizu et al. [14].
@@
Outlook—The physical picture emerging from our work is that LVO is close to a Hund-assisted orbital selective Mott state. Recent experiments have indeed revealed that an insulating state can be induced by Li-intercalation [54] or, together with charge ordering, when subjecting this compound to uniaxial strain [55]. While some qualitative aspects of this physical picture were anticipated before [12, 14, 5], our work demonstrates how key progress on DMFT solvers now allows us to cover the full range of energy scales in this particularly challenging case and reliably establish the physical mechanism at hand. Our work also points at two necessary further developments: (i) including spatial fluctuations and computing spin and charge responses beyond single-site DMFT and (ii) exploring instabilities toward charge and spin ordering dependent on external perturbations such as pressure, uniaxial strain, and chemical substitutions.
@@ Acknowledgements.
Acknowledgements—We acknowledge useful discussions and communications with Ryotaro Arita, Sophie Beck, Riccardo Comin, Andrea Damascelli, Luca de’ Medici, Dongjin Oh, Dennis Huang, Xiangyu Luo, Giorgio Sangiovanni, Hidenori Takagi, Manish Verma, and Jan von Delft. M.G. and U.S. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy-426 EXC-2111-390814868. M.G. and U.S. are grateful to the Flatiron Institute and A.G. to the University of Munich for their hospitality. The Flatiron Institute is a division of the Simons Foundation.

### v2

@@ Abstract
We show that the remarkably small Fermi-liquid coherence scale and large effective mass observed in LiV2O4 are due to the proximity of a Hund-assisted orbital-selective Mott phase (OSMP). Our work is based on an ab initio dynamical mean-field approach, combining several quantum impurity solvers to capture the physics from high to very low temperature. We find that the Hund coupling plays a crucial role in rearranging the orbital populations and in generating the heavy mass and low coherence scale. The latter is found to be approximately 1–2 Kelvin, even though the most correlated orbital is found to be significantly doped ($\sim 10\%$) away from half-filling. A flat quasiparticle band appears near the Fermi level as a result of the strong electronic correlations. Finally, we discuss our results in comparison to experiments.
@@
Figure 1: (a) DFT band structure of LVO for the four $a_{1g}$- and eight $e_{g}^{\pi}$-like bands around the Fermi level. The eigenvalues of the maximally localized Wannier functions (MLWFs) perfectly match the DFT bands [6]. Color: orbital character. Inset: Brillouin zone with high-symmetry points. (b) DFT DOS and projections onto the MLWFs.
@@
We establish a connection to an orbital-selective Mott phase (OSMP) [19, 20, 21, 14, 5, 22], which is crucial in many systems, including iron-based superconductors [23, 24, 25, 26, 27], and show that LVO is close to a Hund-assisted OSMP. Here, the inter-orbital Hund coupling $J$ (and therefore the $e_{g}^{\pi}$ bands) plays a central role. First, it induces a redistribution of the ($a_{1g}$, $e_{g}^{\pi}$) orbital populations per vanadium due to interaction from $(0.4,2\times 0.55)$ in DFT to $(0.9,2\times 0.3)$ in DMFT. Second, it is responsible for the heavy mass and low $T_{\mathrm{FL}}\,$ via the spin-blocking mechanism characteristic of Hund systems in which electrons can only hop between atomic configurations with maximum spin [28, 29, 30, 31]. LVO is shown to be close to an OSMP in which the $a_{1g}$ orbital is prevented to fully localize at $T=0$ due to inter-orbital hopping [32]. Remarkably, we obtain a low $T_{\mathrm{FL}}\,$ already at $\sim\!10\%$ doping of the $a_{1g}$ band, in contrast to the $\sim\!2\%$ doping of Ref. [12].
@@
Figure 2: (a) $\mathrm{Im}\Sigma(i\omega_{n})$ for the $a_{1g}$ orbital (circles) and $e_{g}^{\pi}$ orbitals (crosses) at various $T$ from DFT+DMFT using QMC, TN, and NRG. Inset: zoom at low frequencies. (b) Scattering rate and (c) spectral function at the Fermi level as a function of $T$. Results at $i0^{+}$ in (b)–(c) for QMC and TN were obtained by fitting a fourth-order polynomial to the lowest eight Matsubara points. The inset in (b) reveals the FL behavior of the self-energy, $-\mathrm{Im}\Sigma(i\pi T)\propto T$ [36], and the inset of panel (c) displays the saturation of the spectral function at low $T$.
@@
Crossover to a heavy FL—The emergence of the low-$T$ heavy FL is illustrated in Fig. 2. Figure 2(a) shows the imaginary part of the self-energy for each orbital as a function of Matsubara frequency for several $T$. The $a_{1g}$ self-energy shows almost no $T$ dependence for $\omega_{n}>0.01\,$eV. Only at very low $T$, a sudden drop at low frequencies occurs (best visible in the inset) for calculations below the onset temperature $T_{\mathrm{onset}}\,$$\sim 12$ K. This is a hallmark of strong correlations and corresponds to a sharp crossover from a high-$T$ incoherent regime with a large scattering rate to a low-$T$ coherent regime. This is further supported by Fig. 2(b), which shows the extrapolated zero-frequency value of the $a_{1g}$ scattering rate $-\mathrm{Im}\Sigma_{a_{1g}}(i0^{+})$ as a function of $T$. The inset of Fig. 2(b) depicts $-\mathrm{Im}\Sigma_{a_{1g}}(i\pi T)$, from which we can infer that the FL is formed at 1–2 K when the plot becomes linear [36]. Hence, we conclude that the $T$ range between 12 K, where the bending of $\mathrm{Im}\Sigma(i\omega_{n})$ becomes visible, and 1 K can be identified as the FL crossover regime. The local spectral function (DOS) at the Fermi level $A_{a_{1g}}(0)$ correspondingly increases, as shown in Fig. 2(c). By contrast, the $e_{g}^{\pi}$ self-energy is smaller and depends more weakly on temperature and frequency.
The FL fully develops only below the FL coherence temperature $T_{\mathrm{FL}}\,$$\sim$ 1–2 K, an order of magnitude smaller than $T_{\mathrm{onset}}\,$. The FL regime can also be characterized by $-\mathrm{Im}\Sigma\!\sim\!\omega^{2}\!+\!\pi^{2}T^{2}$ (see End Matter for $T\!=\!0$) for the retarded self-energy and $\chi^{\prime\prime}\!\sim\!\omega$ for the imaginary part of the retarded local spin susceptibility (see End Matter). The quasiparticle weights (from NRG at $T\!=\!0$) are $Z_{a_{1g}}\!\approx\!0.003$ and $Z_{e_{g}^{\pi}}\!\approx\!0.03$.
@@
To analyze the possible proximity of a Mott-insulating state, we perform a numerical experiment in which we add a crystal field to the Wannier Hamiltonian, while keeping the total $d$-electron count equal to 1.5 by adjusting the chemical potential. Figure 4(d) shows $T_{\mathrm{FL}}\,$determined by NRG as a function of the $a_{1g}$ orbital occupancy (lower horizontal scale) or the applied crystal field (upper scale). We see that $T_{\mathrm{FL}}\,$is driven to remarkably low values as half-filling of the $a_{1g}$ orbital is approached, with values as low as $10^{-6}$ K at $1\%$ doping. This can be interpreted as the system reaching an OSMP [19, 20, 21, 14, 5] in which the $e_{g}^{\pi}$ electrons remain itinerant while the $a_{1g}$ ones basically localize, full localization being prevented at $T\!=\!0$ by inter-orbital hopping [32]. The relevance of the OSM phenomenon and of Hund physics as a possible cause for $d$-electron HF behavior was recently emphasized by Crispino et al. [5] and also previously mentioned qualitatively for LVO in the NMR study of Shimizu et al. [14].
@@
Outlook—The physical picture emerging from our work is that LVO is close to a Hund-assisted OSMP. Recent experiments have indeed revealed that an insulating state can be induced by Li-intercalation [61] or, together with charge ordering, when subjecting this compound to uniaxial strain [62]. While some qualitative aspects of this physical picture were anticipated before [12, 14, 5], our work demonstrates how key progress on DMFT solvers now allows us to cover the full range of energy scales in this particularly challenging case and reliably establish the physical mechanism at hand. Our work also points at two necessary further developments: (i) including spatial fluctuations and computing spin and charge responses beyond single-site DMFT and (ii) exploring instabilities toward charge and spin ordering dependent on external perturbations such as pressure, uniaxial strain, and chemical substitutions.
Note added. After completion of this paper, we became aware of the work subsequently posted on arXiv by Backes et al. which also addresses the physics of LiV2O4, with similar physical conclusions [63].
@@ Acknowledgements.
Acknowledgements. We acknowledge useful discussions and communications with Ryotaro Arita, Sophie Beck, Riccardo Comin, Andrea Damascelli, Luca de’ Medici, Dongjin Oh, Dennis Huang, Xiangyu Luo, Giorgio Sangiovanni, Hidenori Takagi, Manish Verma, and Jan von Delft. M.G. and U.S. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2111 No. 390814868. F.B.K. acknowledges funding from the Ministerium für Kultur und Wissenschaft des Landes Nordrhein-Westfalen (NRW Rückkehrprogramm). M.G. and U.S. are grateful to the Flatiron Institute and A.G. to the University of Munich for their hospitality. The Flatiron Institute is a division of the Simons Foundation. Data availability. The data that support the findings of this 337 article are not publicly available. The data are available from 338 the authors upon reasonable request.